OpenStudy (anonymous):

can anyone find the center and the radius for x^2+y^2-6x+2y+9=0

OpenStudy (anonymous):

(x-3)^2+(y+1)^2=1 center (3,-1) radius 1

OpenStudy (anonymous):

how did you actually work that or is it a long process im so confused

OpenStudy (anonymous):

You need to get it into the standard form (x-h)^2 +(y-k)^2=r^2 I grouped the x-terms and y-terms together individually x^2-6x +y^2+2y +9=0 it is called completing the square. x^2-6x since the x^2 coefficient is 1, I take 1/2 of the b coefficient and square it (1/2)(6)=3. 3^2 = 9 so I need x^2-6x+9 so I used the +9 from the problem (x^2-6x+9) +(y^2+2y)=0 (x-3)^2 +(y^2+2y ) =0 now I do the same with y^2+2y (1/2)(2) = 1, 1^2 =1 so I add 1 to both sides of the equation (x-3)^2 +(y^2+2y +1 ) =0+1 (x-3)^2 + (y+1)^2 = 1 1=1^2 so the center is (3,-1) with r=1

OpenStudy (anonymous):

You make it so much easier than my math professor

OpenStudy (anonymous):

glad I could help

OpenStudy (anonymous):

:)